The nimbus package^{∗}

Peter Dyballa

1997/01/23

1 Introduction

This file contains the external font information needed to load the Nimbus fonts, and a suitable package file to utilize them.

From this file all .fd files (font definition files) for the PostScript fonts with old encoding (OT1) are created, as well as those with the new encoding (T1).

2 Customization

The usual customization will be to add support for additional combinations in this file, such as a derived bold slanted font.

3 The docstrip modules

As always we begin by identifying the latest version of the files in the log file. The explicit spaces are necessary in an .fd file and the \string guards against situations where ` is active.

1 ⟨∗package⟩

 2 \ProvidesPackage{nimbus}[\filedate\space\fileversion\space
 3 PSNFSS2e LaTeX2e package]
 4 \renewcommand{\sfdefault}{unms}
 5 \renewcommand{\rmdefault}{unm}
 6 \def\bfdefault{b}

7 ⟨/package⟩

 8 % \subsection{EC-encoded URW NimbusRoman-Regular}
 9 %

10 ⟨∗T1⟩

11 \typeout{File\space t1unm.fd\space loading \space URW\space NimbusRoman-Regular}%
12 \DeclareFontFamily{T1}{unm}{}%
13 \DeclareFontShape{T1}{unm}{m}{n}{<->unm8t}{}%
14 \DeclareFontShape{T1}{unm}{m}{sc}{<->unmc8t}{}%

∗This file has version number 3, last revised 1997/01/23.

15 \DeclareFontShape{T1}{unm}{m}{it}{<->sub * unm/m/sl}{}%
16 \DeclareFontShape{T1}{unm}{m}{sl}{<->unmo8t}{}%
17 \DeclareFontShape{T1}{unm}{mn}{n}{<->unmn8t}{}%
18 \DeclareFontShape{T1}{unm}{b}{n}{<->sub * ptmr/b/n}{}%
19 \DeclareFontShape{T1}{unm}{b}{it}{<->sub * ptm/b/sl}{}%
20 \DeclareFontShape{T1}{unm}{b}{sc}{<->sub * ptm/b/sc}{}%
21 \DeclareFontShape{T1}{unm}{b}{sl}{<->sub * ptm/b/sl}{}%

22 ⟨/T1⟩ 23 %

24 % \subsection{EC-encoded URW NimbusSans-Regular}

25 % 26 ⟨∗T1sans⟩

27 \typeout{File\space t1unms.fd\space loading \space URW\space NimbusSans-Regular}%
28 \DeclareFontFamily{T1}{unms}{}%
29 \DeclareFontShape{T1}{unms}{m}{n}{<->unms8t}{}%
30 \DeclareFontShape{T1}{unms}{m}{sc}{<->unmsc8t}{}%
31 \DeclareFontShape{T1}{unms}{m}{it}{<->sub * unms/m/sl}{}%
32 \DeclareFontShape{T1}{unms}{m}{sl}{<->unmso8t}{}%
33 \DeclareFontShape{T1}{unms}{mn}{n}{<->unmsn8t}{}%
34 \DeclareFontShape{T1}{unms}{b}{n}{<->sub * phv/b/n}{}%
35 \DeclareFontShape{T1}{unms}{b}{it}{<->sub * phv/b/sl}{}%
36 \DeclareFontShape{T1}{unms}{b}{sc}{<->sub * phv/b/sc}{}%
37 \DeclareFontShape{T1}{unms}{b}{sl}{<->sub * phv/b/sl}{}%

38 ⟨/T1sans⟩ 39 %

40 % \subsection{Old \TeX-encoded URW NimbusRoman-Regular}

41 % 42 ⟨∗OT1⟩

43 \typeout{File\space ot1unm.fd\space loading \space URW\space NimbusRoman-Regular}%
44 \DeclareFontFamily{OT1}{unm}{}%
45 \DeclareFontShape{OT1}{unm}{m}{n}{<->unm7t}{}%
46 \DeclareFontShape{OT1}{unm}{m}{sc}{<->unmc7t}{}%
47 \DeclareFontShape{OT1}{unm}{m}{it}{<->sub * unm/m/sl}{}%
48 \DeclareFontShape{OT1}{unm}{m}{sl}{<->unmo7t}{}%
49 \DeclareFontShape{OT1}{unm}{mn}{n}{<->unmn7t}{}%
50 \DeclareFontShape{OT1}{unm}{b}{n}{<->sub * ptm/b/n}{}%
51 \DeclareFontShape{OT1}{unm}{b}{it}{<->sub * ptm/b/sl}{}%
52 \DeclareFontShape{OT1}{unm}{b}{sc}{<->sub * ptm/b/sc}{}%
53 \DeclareFontShape{OT1}{unm}{b}{sl}{<->sub * ptm/b/sl}{}%

54 ⟨/OT1⟩ 55 %

56 % \subsection{Old \TeX-encoded URW NimbusSans-Regular}

57 % 58 ⟨∗OT1sans⟩

59 \typeout{File\space ot1unms.fd\space loading \space URW\space NimbusSans-Regular}%
60 \DeclareFontFamily{OT1}{unms}{}%
61 \DeclareFontShape{OT1}{unms}{m}{n}{<->unms7t}{}%
62 \DeclareFontShape{OT1}{unms}{m}{sc}{<->unmsc7t}{}%
63 \DeclareFontShape{OT1}{unms}{m}{it}{<->sub * unms/m/sl}{}%
64 \DeclareFontShape{OT1}{unms}{m}{sl}{<->unmso7t}{}%

65 \DeclareFontShape{OT1}{unms}{mn}{n}{<->unmsn7t}{}%
66 \DeclareFontShape{OT1}{unms}{b}{n}{<->sub * phv/b/n}{}%
67 \DeclareFontShape{OT1}{unms}{b}{it}{<->sub * phv/b/sl}{}%
68 \DeclareFontShape{OT1}{unms}{b}{sc}{<->sub * phv/b/sc}{}%
69 \DeclareFontShape{OT1}{unms}{b}{sl}{<->sub * phv/b/sl}{}%

70 ⟨/OT1sans⟩ 71 %

Index

Numbers written in italic refer to the page where the corresponding entry is described, the ones underlined to the code line of the definition, the rest to the code lines where the entry is used.

B

\bfdefault . . . . . . . . 6

D

\DeclareFontFamily .

…. 12, 28, 44, 60

\DeclareFontShape .

… 13–21, 29– 37, 45–53, 61–69

\def . . . . . . . . . . . . . . 6

F \filedate . . . . . . . . . 2

                          \fileversion . . . . . . . 2

P

                          \ProvidesPackage . . . 2

R

                          \renewcommand . . . . 4, 5
                          \rmdefault . . . . . . . . 5

S

                                                   \sfdefault . . . . . . . . 4

\space . 2, 11, 27, 43, 59

                                                   \subsection 8, 24, 40, 56

T \TeX . . . . . . . . . . . 40, 56 \typeout . 11, 27, 43, 59